# Dynamical activity universally bounds precision of response

### v2

@@ Introduction
where $\mathcal{A}$ is the DA [7, 8, 9, 10, 21, 11] The TUR (1) and the KUR (2) have complementary roles and they can be derived from a unified thermodynamic-kinetic uncertainty relation (TKUR) [6].
Figure 1: Monte Carlo simulation of a time-integrated current during the relaxation process of a three-state model. We use blue color to depict the process with a certain value of the control parameter $\theta$ and use red color for a perturbed control parameter $\theta+{\rm d}\theta$. The two thick curves represent the average current and each thin curve is a stochastic trajectory. The deviation of the average values represents ${\rm d}J$ due to a change in ${\rm d}\theta$.
@@
Concurrently, the development of finding fundamental principles characterizing nonequilibrium systems’ static response to external perturbations is also prosperous [24, 25, 26, 27, 28, 29, 30] (see Fig. 1 for a schematic illustration), apart from the seminal fluctuation-dissipation relation [31, 32] and its generalizations in the far-from-equilibrium regime [33, 34, 35, 36, 37, 8, 38, 39, 40, 41]. Notably, Ptaszyński et al. [42] very recently conjectured a novel metric that assesses the precision of the current response to kinetic perturbations and conjectured a response thermodynamic uncertainty relation (R-TUR) valid for the NESS of a homogeneous Markov jump process, expressed as:
@@
where $p_{i}(t)$ is the probability of state $i$ at time $t$, $W_{ij}^{\nu}(t)$ is the time-dependent transition rate from state $j$ to state $i$ via channel $\nu$ (e.g., by coupling with the $\nu$th reservoir), and $W_{ij}(t)=\sum_{\nu}W_{ij}^{\nu}(t)$. The transition rate has a general form as $W_{ij}^{\nu}=\exp(B_{ij}^{\nu}+E_{j}^{\nu}+S_{ij}^{\nu}/2)$, where $B_{ij}^{\nu}$ is the symmetric part, $E_{j}^{\nu}$ is the vertex part, and $S_{ij}^{\nu}$ is the anti-symmetric contribution [25]. Each of these three contributions can be controlled independently. Different expressions of them correspond to different types of transition rates, such as the Arrhenius type [47], the phenotypic transition rates [48], and the phonon spectral type [49]. We generically parameterize the transition rates as $W_{ij}^{\nu}(\theta,t)$, where $\theta$ is the control parameter. The only requirement is that the control parameter $\theta$ does not explicitly depend on the time $t$.
@@
NESS Relaxation Time-dependent driving Observable Saturable for discrete-state
TUR [4, 5] (1) yes no no current no
KUR [18, 19, 20, 17] (2) yes yes no generic no
R-TUR [42] (3) conjectured unknown unknown current unknown
R-KUR (6) yes yes yes generic yes
Table 1: Comparison of the range of applicability for the TUR, the KUR, the R-TUR, and our R-KUR.
By comparing inequalities (3) and (6), the R-KUR can be regarded as a complementary counterpart of the R-TUR in terms of the DA, yet transcending its limitation on perturbations of the kinetic barrier or time-independent transition rates. In fact, the R-KUR can be applied to arbitrary Markov jump processes as summarized in Table 1 and does not necessitate perturbations to the driving speed or observation time, making it more versatile in applications. We also note that the R-KUR is valid for any process that can be described by the master equation (5), without being restricted to thermodynamic systems. Thermodynamic interpretation can be attained by introducing local detailed balance conditions for transition rates, namely $W_{ij}^{\nu}\exp(-\beta_{\nu}\epsilon_{j})=W_{ji}^{\nu}\exp(-\beta_{\nu}\epsilon_{i})$ with $\beta_{\nu}$ being the inverse temperature of the $\nu$th heat bath and $\epsilon_{i}$ the energy of state $i$ [50].
@@
Suppose that a random walker jumps from the position $i$ to the position $i\pm 1$ at a Arrhenius-type rate $k_{\pm}:=\exp{(\theta\pm\alpha/2)}$ [47], with $\theta>0$ being the symmetric Kramer’s rate [60] controlled by the potential barrier height, and the anti-symmetric part $\alpha>0$ determined by an external potential gradient. The observable is chosen as the current $J$ that flows from the negative direction to the positive direction in the NESS. We first introduce a perturbation in the control parameter $\theta$ while keeping $\alpha$ fixed. As calculated in Methods, we find that the R-TUR (3) conjectured in Ref. [42] becomes exactly the same as the inequality (4) in the original paper of the TUR [4]. It is explicitly written as
@@
which requires $\alpha>2.399$. Therefore, the R-TUR is possibly broken while $2.3990<\alpha<$ 2.399.
@@
Figure 2: (a) Schematic of a two-level system $S$ coupled with a cold heat bath and a hot bath whose inverse temperatures are denoted as $\beta_{c}$ and $\beta_{h}$, respectively. The energy gap between the 2 levels 0 and 1 is $\Delta$ which can be controlled externally. The transitions can be caused by coupling with either of the baths. (b) Efficiency of the bound $\eta$ versus the dissipation-activity ratio $\sigma/\mathcal{A}$. Red: the time-integrated current flowing into the cold bath; Blue: the total number of transitions induced by coupling to the cold bath. The bound tends to saturation with a large dissipation-activity ratio, indicating that the tightness of the bound increases as the system increasingly deviates from the equilibrium.
@@
The second example involves a two-level system in contact with two baths at different temperatures. As schematically depicted in Fig. 2(a), the energy gap is denoted by $\Delta$. We assume the local detailed balance condition $W^{\nu}_{01}/W^{\nu}_{10}=e^{\beta_{\nu}\Delta}$ for each bath $\nu(=c,h)$ with inverse temperature $\beta_{\nu}$. Suppose that external control can manipulate the energy gap $\Delta$ as a function of time $t$
@@
We focus on two key observables: one is the time-integrated current flowing into the cold bath $J_{\rm e}[\omega]:=\sum_{i\neq j}(n_{ij}^{c}[\omega]-n_{ji}^{c}[\omega])$; the other is the total number of transitions, or the traffic, induced by coupling to the cold bath $A_{\rm e}^{c}[\omega]:=\sum_{i\neq j}n_{ij}^{c}[\omega]$, where $n_{ij}^{c}[\omega]$ counts the total number of transitions from state $j$ to state $i$ via channel $c$. Their average and variance are calculated by full counting statistics (FCS). See Methods for details. In the numerical simulation, we set $W_{10}^{c}=W_{10}^{h}=1$, $\beta_{h}=1$. We randomly sampled $\beta_{c},\varepsilon,C,s,a$, $\omega$ and the time duration $\tau$. Figure 2(b) shows a plot of $10^{5}$ data points for each observable, red for the current and blue for the counting observable, with the efficiency of the bound $\eta:={(\partial_{\theta}{O})^{2}}/{\langle\langle{O}\rangle\rangle}a_{\max}^{2}\mathcal{A}\leq 1$ on the y-axis and the dissipation-activity ratio $\sigma/\mathcal{A}$ on the x-axis.
@@
It might be natural to imagine a unified R-TKUR that sets a bound on the precision of response in terms of both the EP and the DA, in analogy to the TKUR derived in Ref. [6]. However, it is highly non-trivial to prove or falsify the existence of such a bound. The difficulty lies in two folds: first, the examples suggest that there may not be a unified bound for generic transition rates because the validation of an entropic bound requires the symmetric assumption; second, the R-TUR so far remains a conjecture and is proved only for the linear response regime and unicycle networks [42]. Nevertheless, Fig. 2(b) implies the possible existence of a tighter bound as a function of the dissipation-activity ratio $\sigma/\mathcal{A}$.
Another feature of our results is that all involved quantities are experimentally measurable. The physical system can be implemented with a single defect center in diamond [73], single-electron boxes at low temperatures [74, 75], molecular motors [76], NMR setups [77], nanoscale electronic conductors [78], superconducting NISQ quantum processors [79], etc, in which stochastic trajectories can be tracked with high resolution.
@@ Methods
The mathematical foundation for deriving our bound is the generalized Cramér-Rao inequality [94]. This inequality addresses scenarios where the probability distribution of a random variable $X$ depends on a parameter $\theta$, aiming to estimate the function $\psi(\theta)$ of $X$ through a function $\Psi(X)$. The inequality is explicitly given by
@@
where $I(\theta)$ is the Fisher information.
@@
$a_{\max}:=\mathop{\max_{i\neq j,\nu}}_{t\in[0,\tau]}|\partial^{\nu}_{\theta}\ln W_{ij}(\theta,t)|,$ (23)
@@
$\mathcal{W}_{\chi}(t)=\begin{pmatrix}-W_{10}^{h}-W_{10}^{c}&W_{01}^{h}+W_{01}^{c}e^{\chi}\\ W_{10}^{h}+W_{10}^{c}e^{-\chi}&-W_{01}^{h}-W_{01}^{c}\end{pmatrix}$ (33)
is the transition rate matrix with a counting field $\chi$ for the first observable, i.e., the empirical accumulated current that flows into the cold bath $J^{c}_{\rm e}[\omega]:=\sum_{i\neq j}(n_{ij}^{c}[\omega]-n_{ji}^{c}[\omega])$. Its average and variance are calculated by the derivatives of the cumulant generating function as
$J^{c}=\left.\frac{\partial G_{\chi}(\tau)}{\partial\chi}\right|_{\chi=0},\quad\langle\langle J^{c}\rangle\rangle=\left.\frac{\partial^{2}G_{\chi}(\tau)}{\partial\chi^{2}}\right|_{\chi=0}.$ (34)
The second observable we simulated is the empirical time-integrated traffic due to coupling with the cold bath, defined as $A_{\rm e}^{c}[\omega]:=\sum_{i\neq j}n_{ij}^{c}[\omega]$. The ensemble average and the variance can also be calculated via FCS by introducing a different counting field $\lambda$ as [99]
@@
$\mathcal{W}_{\lambda}(t)=\begin{pmatrix}-W_{10}^{h}-W_{10}^{c}&W_{01}^{h}+W_{01}^{c}e^{\lambda}\\ W_{10}^{h}+W_{10}^{c}e^{\lambda}&-W_{01}^{h}-W_{01}^{c}\end{pmatrix}.$ (36)
In the numerical simulation of both observables, we sample each of the parameters contained in the energy gap $\Delta$ uniformly from the following intervals: $\beta_{c}\in[1,5]$, $\varepsilon\in[0.1,1]$, $C\in[0,2]$, $s\in[-0.5,0.5]$, $a\in[0,1]$, $\omega\in[0,5]$, and $\tau\in[0,1]$.
@@
The Fisher information for a feedback-controlled process takes the following form [65] {strip}

### v3

@@ Introduction
where $\mathcal{A}$ is the DA [7, 8, 9, 10, 21, 11]. The TUR (1) and the KUR (2) have complementary roles and they can be derived from a unified thermodynamic-kinetic uncertainty relation (TKUR) [6].
Figure 1: Monte Carlo simulation of a time-integrated current $J$ during the relaxation process of time $t$ of a three-state model. We use blue color to depict the process with a certain value of the control parameter $\theta$ and use red color for a perturbed control parameter $\theta+{\rm d}\theta$. The two thick curves represent the average current and each thin curve is a stochastic trajectory. The deviation of the average values represents ${\rm d}J$ due to a change ${\rm d}\theta$.
NESS Relaxation Time-dependent driving Observable Equality attainability for discrete-state
TUR [4, 5] (1) yes no no current no
KUR [18, 19, 20, 17] (2) yes yes no generic no
R-TUR [22, 23] (3) yes unknown unknown current unknown
R-KUR (7) yes yes yes generic yes
Table 1: Comparison of the range of applicability for the TUR, the KUR, the R-TUR, and our R-KUR. The abbreviation NESS means the nonequilibrium steady state.
@@
Concurrently, the development of finding fundamental principles characterizing nonequilibrium systems’ static response to external perturbations is also prosperous [26, 27, 28, 29, 30, 31, 32] (see Fig. 1 for a schematic illustration), apart from the seminal fluctuation-dissipation relation [33, 34] and its generalizations in the far-from-equilibrium regime [35, 36, 37, 38, 39, 8, 40, 41, 42, 43]. Notably, a response thermodynamic uncertainty relation (R-TUR) valid for the NESS of a homogeneous Markov jump process have been proposed very recently [22, 23], expressed as:
@@
where $p_{i}(t)$ is the probability of state $i$ at time $t$, $W_{ij}^{\nu}(t)$, for $i\neq j$, is the time-dependent transition rate from state $j$ to state $i$ via channel $\nu$ (e.g., by coupling with the $\nu$th reservoir), and $W_{ij}(t)=\sum_{\nu}W_{ij}^{\nu}(t)$. By conservation of probability, the escape rate of leaving a state $i$ is then $W_{ii}=-\sum_{j(\neq i)}W_{ij}$. With the master equation, the total EP for an interval $[0,\tau]$ is explicitly given by [48, 49]
$\sigma=\int_{0}^{\tau}{\rm d}t\sum_{i\neq j,\nu}W_{ij}^{\nu}(\theta,t)p_{j}(t)\ln\frac{W_{ij}^{\nu}(\theta,t)p_{j}(t)}{W_{ji}^{\nu}(\theta,t)p_{i}(t)}.$ (6)
The transition rate has a general form as $W_{ij}^{\nu}=\exp(B_{ij}^{\nu}+E_{j}^{\nu}+S_{ij}^{\nu}/2)$, where $B_{ij}^{\nu}$ is the symmetric part, $E_{j}^{\nu}$ is the vertex part, and $S_{ij}^{\nu}$ is the anti-symmetric contribution [27]. Each of these three contributions can be controlled independently. Different expressions of them correspond to different types of transition rates, such as the Arrhenius type [50], the phenotypic transition rates [51], and the phonon spectral type [52]. We generically parameterize the transition rates as $W_{ij}^{\nu}(\theta,t)$, where $\theta$ is the control parameter. The only requirement is that the control parameter $\theta$ does not explicitly depend on the time $t$.
@@
By comparing inequalities (3) and (7), the R-KUR can be regarded as a complementary counterpart of the R-TUR in terms of the DA, yet transcending its limitation on perturbations of the kinetic barrier or time-independent transition rates. In fact, the R-KUR can be applied to arbitrary Markov jump processes as summarized in Table 1 and does not necessitate perturbations to the driving speed or observation time, making it more versatile in applications. We also note that the R-KUR is valid for any process that can be described by the master equation (5), without being restricted to thermodynamic systems. Thermodynamic interpretation can be attained by introducing local detailed balance conditions for transition rates, namely $W_{ij}^{\nu}\exp(-\beta_{\nu}\omega_{j}^{\nu})=W_{ji}^{\nu}\exp(-\beta_{\nu}\omega_{i}^{\nu})$, where $\beta_{\nu}$ is the inverse temperature of the $\nu$th reservoir and $\omega_{i}^{\nu}:=\epsilon_{i}-\mu^{\nu}N_{i}$ is the grand potential associated with the $\nu$th reservoir, with $\epsilon_{i}$ being the energy level of the $i$th inner state of the system, $\mu^{\nu}$ the chemical potential of the $\nu$th reservoir, and $N_{i}$ the number of particles in state $i$ of the system [53, 54].
Figure 2: Numerical simulation of the efficiency of the bound $\eta$ of the R-TURs and the R-KURs as a function of the dissipation-activity ratio $\sigma/\mathcal{A}$ for Example 1. Green, blue, red, and yellow curves represent the efficiency calculated by inequalities (11)-(14) about the R-TUR and the R-KUR with perturbation in either the symmetric parameter $\theta$ or the anti-symmetric parameter $\alpha$, respectively.
@@
Suppose that a random walker jumps from the position $i$ to the position $i\pm 1$ at an Arrhenius-type rate $k_{\pm}:=\exp{(\theta\pm\alpha/2)}$ [50], with $\theta>0$ being the symmetric Kramer’s rate [64] controlled by the potential barrier height, and the anti-symmetric part $\alpha>0$ determined by an external potential gradient. The observable is chosen as the current $J$ that flows from the negative direction to the positive direction in the NESS. We first introduce a perturbation in the control parameter $\theta$ while keeping $\alpha$ fixed. As calculated in Methods, we find that the R-TUR (3) becomes exactly the same as the inequality (4) in the original paper of the TUR [4]. It is explicitly written as
@@
which requires $\alpha>2.399$. Therefore, the R-TUR is possibly broken while $2.3990<\alpha<$ 2.399. We perform numerical simulations for the above 4 inequalities in Figure 2 by sampling $\theta$, $\alpha$, and $\tau\in[0,10]$ for $10^{5}$ data points. We define the efficiency of the bound $\eta:={(\partial_{\theta}{O})^{2}}/{\langle\langle{O}\rangle\rangle}a_{\max}^{2}\mathcal{A}$ for the R-KUR and $\eta:=2{(\partial_{\theta}{O})^{2}}/{\langle\langle{O}\rangle\rangle}a_{\max}^{2}\sigma$ for the R-TUR. The validity of the bound implies $\eta\leq 1$; the closer $\eta$ is to 1, the tighter the bound is.
@@
Figure 3: Numerical illustration with a two-level system coupled to two heat baths. (a) Schematic of a two-level system $S$ coupled with a cold heat bath and a hot bath whose inverse temperatures are denoted as $\beta_{\rm c}$ and $\beta_{\rm h}$, respectively. The energy gap between the 2 levels 0 and 1 is $\Delta$ which can be controlled externally. The transitions can be caused by coupling with either of the baths with 4 transition rates $W_{ij}^{\rm h(c)}$, $i,j\in\{0,1\}$. (b) Efficiency of the bound $\eta$ versus the dissipation-activity ratio $\sigma/\mathcal{A}$. Red dots are plotted with a time-integrated current $J^{\rm c}$ flowing into the cold bath; blue dots are plotted with the total number of transitions induced by coupling to the cold bath. The bound tends to saturation with a large dissipation-activity ratio, indicating that the tightness of the bound increases as the system increasingly deviates from the equilibrium.
@@
The second example involves a two-level system in contact with two baths at different temperatures. As schematically depicted in Fig. 3(a), the energy gap is denoted by $\Delta$. We assume the local detailed balance condition $W^{\nu}_{01}/W^{\nu}_{10}=e^{\beta_{\nu}\Delta}$ for each bath $\nu(={\rm c,h})$ with inverse temperature $\beta_{\nu}$. Suppose that external control can manipulate the energy gap $\Delta$ as a function of time $t$
@@
We focus on two key observables: one is the time-integrated current flowing into the cold bath $J_{\rm e}[\omega]:=\sum_{i\neq j}(n_{ij}^{\rm c}[\omega]-n_{ji}^{\rm c}[\omega])$; the other is the total number of transitions, or the traffic, induced by coupling to the cold bath $A_{\rm e}^{\rm c}[\omega]:=\sum_{i\neq j}n_{ij}^{\rm c}[\omega]$, where $n_{ij}^{\rm c}[\omega]$ counts the total number of transitions from state $j$ to state $i$ via channel c. Their average and variance are calculated by full counting statistics (FCS). See Methods for details. In the numerical simulation, we set $W_{10}^{\rm c}=W_{10}^{\rm h}=1$, $\beta_{\rm h}=1$. We randomly sampled $\beta_{\rm c},\varepsilon,C,s,a$, $\omega$ and the time duration $\tau$. Figure 3(b) shows a plot of $10^{5}$ data points for each observable, red for the current and blue for the counting observable, with the efficiency of the bound $\eta:={(\partial_{\theta}{O})^{2}}/{\langle\langle{O}\rangle\rangle}a_{\max}^{2}\mathcal{A}\leq 1$ on the y-axis and the dissipation-activity ratio $\sigma/\mathcal{A}$ on the x-axis.
@@
Equality condition
Example 1 has demonstrated that the R-KUR can attain its equality. Here we present another sufficient condition of equality. Consider the steady state of a homogeneous Markov jump process with only one channel. Inspired by Ref. [65], we focus on the response around $\theta=0$. The parametrization takes a special form as $W_{ij}(\theta)=W_{ij}\exp(\theta Z_{ij})$ for $i\neq j$ and $W_{jj}(\theta)=-\sum_{i}W_{ij}\exp(\theta Z_{ij})$, where $Z_{ij}=(W_{ij}p_{j}^{\text{ss}}-W_{ji}p_{i}^{\text{ss}})/(W_{ij}p_{j}^{\text{ss}}+W_{ji}p_{i}^{\text{ss}})$ (the superscript ss denotes steady-states) and $W_{ij}$’s are unperturbed transition rates. The empirical observable should take the following form:
$O[\omega]$ $=\mathcal{J}+\mathcal{I}$ (16)
$:=\sum_{i\neq j}d_{ij}(n_{ij}-n_{ji})+\sum_{i}q_{i}\int_{0}^{\tau}\text{d}t\delta_{w(t),i},$
where $d_{ij}:=(W_{ij}p^{\text{ss}}_{j}-W_{ji}p^{\text{ss}}_{i})/(W_{ij}p^{\text{ss}}_{j}+W_{ji}p^{\text{ss}}_{i})$ and $q_{j}:=-\sum_{i}W_{ij}(W_{ij}p^{\text{ss}}_{j}-W_{ji}p^{\text{ss}}_{i})/(W_{ij}p^{\text{ss}}_{j}+W_{ji}p^{\text{ss}}_{i})$. The first term $\mathcal{J}$ is the time-integrated current where $n_{ij}$ is the number of jumps from state $j$ to $i$. The second term $\mathcal{I}$ is the so-called time-integrated empirical measure where $w(t)$ is the state of the system at time $t$. It is proved that the average of $O$ vanishes [65].
The setting above saturates the Cauchy-Schwarz inequality (23) and the generalized Cramér-Rao inequality since $O[\omega]=\partial_{\theta}\ln P_{\theta}[\omega]$ [65] (see Methods for definitions of the symbols). To attain the equality in inequality (25), one further needs the condition that $|\partial_{\theta}\ln W_{ij}(\theta)|$ is a constant, i.e., $|Z_{ij}|$ is a constant, which is always valid for two-state systems and can be satisfied for systems with more states.
@@ Conclusions
Another feature of our results is that all involved quantities are experimentally measurable. Quantities involved in various results derived in stochastic thermodynamics, such as EP, DA, and the current statistics, can be measured from the statistics of stochastic jumps [77, 78, 79, 80]. It is worth mentioning that the measurement of DA can sometimes be more accessible than that of EP, as the latter requires distinguishing the direction of jumps. Candidates for physical implementation that allow stochastic trajectories to be tracked with a high resolution include a single defect center in diamond [81], single-electron boxes at low temperatures [82, 83], molecular motors [84], NMR setups [85], nanoscale electronic conductors [86], and superconducting NISQ quantum processors [87], etc.
@@
Note added: we have noticed Ref. [101], which was uploaded to the preprint server subsequent to our submission. Part of its main results is congruent with applying our R-KUR to steady states. This attests to the solidity of our work and indicates that the two studies are mutually complementary.
@@ Methods
The mathematical foundation for deriving our bound is the generalized Cramér-Rao inequality [102]. This inequality addresses scenarios where the probability distribution of a random variable $X$ depends on a parameter $\theta$, aiming to estimate the function $\psi(\theta)$ of $X$ through an estimator $\Psi(X)$. Hence, $\langle\Psi(X)\rangle_{\theta}:=\int_{X}{\rm d}X\Psi(X)P_{\theta}(X)=\psi(\theta)$. The generalized Cramér-Rao bound is explicitly given by
@@
where $I(\theta)$ is the Fisher information, and can be straightforwardly proved with the Cauchy-Schwarz inequality as follows:
$\langle\langle\Psi\rangle\rangle_{\theta}I(\theta)$ $=\int\text{d}X(\Psi-\psi)^{2}P_{\theta}\int\text{d}X(\partial_{\theta}\ln P_{\theta})^{2}P_{\theta}$ (23)
$\geq\left[\int\text{d}X(\Psi-\psi)(\partial_{\theta}\ln P_{\theta})P_{\theta}\right]^{2}$
$=[\text{d}_{\theta}\psi(\theta)]^{2}.$
@@
$a_{\max}:=\mathop{\max_{i\neq j,\nu}}_{t\in[0,\tau]}|\partial_{\theta}\ln W_{ij}^{\nu}(\theta,t)|,$ (26)
@@
$\mathcal{W}_{\chi}(t)=\begin{pmatrix}-W_{10}^{\rm h}-W_{10}^{\rm c}&W_{01}^{\rm h}+W_{01}^{\rm c}e^{\chi}\\ W_{10}^{\rm h}+W_{10}^{\rm c}e^{-\chi}&-W_{01}^{\rm h}-W_{01}^{\rm c}\end{pmatrix}$ (36)
is the transition rate matrix with a counting field $\chi$ for the first observable, i.e., the empirical accumulated current that flows into the cold bath $J^{\rm c}_{\rm e}[\omega]:=\sum_{i\neq j}(n_{ij}^{\rm c}[\omega]-n_{ji}^{\rm c}[\omega])$. Its average and variance are calculated by the derivatives of the cumulant generating function as
$J^{\rm c}=\left.\frac{\partial G_{\chi}(\tau)}{\partial\chi}\right|_{\chi=0},\quad\langle\langle J^{\rm c}\rangle\rangle=\left.\frac{\partial^{2}G_{\chi}(\tau)}{\partial\chi^{2}}\right|_{\chi=0}.$ (37)
The second observable we simulated is the empirical time-integrated traffic due to coupling with the cold bath, defined as $A_{\rm e}^{\rm c}[\omega]:=\sum_{i\neq j}n_{ij}^{\rm c}[\omega]$. The ensemble average and the variance can also be calculated via FCS by introducing a different counting field $\lambda$ as [106]
@@
$\mathcal{W}_{\lambda}(t)=\begin{pmatrix}-W_{10}^{\rm h}-W_{10}^{\rm c}&W_{01}^{\rm h}+W_{01}^{\rm c}e^{\lambda}\\ W_{10}^{\rm h}+W_{10}^{\rm c}e^{\lambda}&-W_{01}^{\rm h}-W_{01}^{\rm c}\end{pmatrix}.$ (39)
In the numerical simulation of both observables, we sample each of the parameters contained in the energy gap $\Delta$ uniformly from the following intervals: $\beta_{\rm c}\in[1,5]$, $\varepsilon\in[0.1,1]$, $C\in[0,2]$, $s\in[-0.5,0.5]$, $a\in[0,1]$, $\omega\in[0,5]$, and $\tau\in[0,1]$.
@@
The Fisher information for a feedback-controlled process takes the following form [49]
@@
Code availability
All the computational codes that were used to generate the data presented in this study are available from the corresponding authors upon reasonable request.